This is a prob/stat question. Hopefully someone will help since I am so lost right now. Trying to prepare for a test.
In a study the carapace lengths (in mm) of Thenus lobster caught near Singapore measured: 78, 66, 65, 63, 60, 60, 58, 56, 52, and 50 (the population data is approximately normal). Find (round to nearest hundredths):
a. The sample mean = __________ (s = 7.97)
b. The 90% confidence intervals for the population mean _________________.
c. The 70% confidence intervals for the population mean _________________.
d. The 80% confidence intervals for the population mean _________________.
e. How many would they need to sample to be 97% sure to be within 0.07 of the population mean? ______________.
f. How many more (than the original 10) would they need to sample to be 64% sure to be within 0.1 of the population mean? ______________.

Answer by stanbon(75887)

This is a prob/stat question. Hopefully someone will help since I am so lost right now. Trying to prepare for a test.
In a study the carapace lengths (in mm) of Thenus lobster caught near Singapore measured: 78, 66, 65, 63, 60, 60, 58, 56, 52, and 50 (the population data is approximately normal). Find (round to nearest hundredths):
a. The sample mean = 60.8____ (s = 7.97)
-------------------
b. The 90% confidence intervals for the population mean _________________.
ME = 1.645 = 4.416
CI::: 60.8-5.42 < u < 60.8+4.42
---------------------------
c. The 70% confidence intervals for the population mean _________________.
Has a different ME:
ME = 1.0364[7.97/sqrt(10)] = 2.61
CI: 60.8-2.61 < u < 60.8+2.61
---------------------------
d. The 80% confidence intervals for the population mean _________________.
Use a different ME
---------------------------------
e. How many would they need to sample to be 97% sure to be within 0.07 of the population mean? ______________.
n = [z*s/E]^2 = [2.17*7.97/0.07]^2 = 61,049 when rounded up
==========================
f. How many more (than the original 10) would they need to sample to be 64% sure to be within 0.1 of the population mean? ______________.
n = [0.9154*7.97/0.01]^2 = 532238 when rounded up
---
You would need 53228 more samples.
=========================
Cheers,
Stan H.
===============